diaz is 1 year more than 3 times as old as his daughter. 4 years from now he will lack 7 years from being 3 times as old as his daughter is then. find each age. this is what i have tried, with the daughter being x she is x now and x plus 4 then he is 1 plus 3x now and 4 plus 1 plus 3x then - so... 1 plus 3x equals x now and 4 plus 1 plus 3x equals x plus 4 then. i think the answer comes out to be 31 and 10, but don't know if this is the right way to get there. am i close?? thank you so much for the help!!
Answer by stanbon(75887)

diaz is 1 year more than 3 times as old as his daughter.
NOW DATA:
Let her daughter's age be "x" yrs. old
Diaz's age is "3x+1" yrs. old.
-------------------------------------
4 yrs. from NOW DATA:
Daughter will be "x+4" yrs.old
Diaz will be 3x+1+4 = "3x+5" yrs old.
-------------------------------------------
4 years from now he will lack 7 years from being 3 times as old as his daughter is then. find each age.
----------------
EQUATION:
3x+5+7 = 3(x+4)
3x+12 = 3x+12
==================
It looks like Diaz's daughter could be any age
and Diaz will be 3 times that + 1 yr.
=============================
Cheers,
Stan H.